I am stuck on a problem. I need to find the value of 'x' and simplify. I've tried 3 different ways to do it and none are coming out right, (according the the back of the book).
The equation is:
2(5x-3)-(2x-4)=5-(6x+1)
please help!!!

Answer by Alan3354(69443)

2(5x-3)-(2x-4)=5-(6x+1)
10x-6 -2x+4 = 5 -6x-1
8x - 2 = 4 - 6x
14x = 6
x = 3/7